LOG-IN
Displaying reports 3261-3280 of 3360.Go to page Start 160 161 162 163 164 165 166 167 End
KAGRA VIS (General)
Print this report.
RyutaroTakahashi - 10:09, Thursday 18 June 2015 (108)Get code to link to this report
Adjustment of breadboard in MCE
Member: R. Takahashi, N. Ohishi, K. Doi

June 11
1) Clean the floor of clean booth.
2) Open the chamber.
3) Adjust the height of breadboard. The height was set to 490mm +/- 1mm from the base plate.
4) Measure the position of the breadboard by the straight edge and the pendulum. The position was deviated by -1.0mm, -1.75mm from the center of the chamber.

June 12
1) Measure the rotation of the breadboard by the straight edge and the pendulum. The position of support bolt was on the Y line within +/- 0.5mm.
2) Pick up some parts from the bottom of chamber. We found M10 long bolt, Al foil, I-bolt, and CCP base.
3) Close the chamber.
KAGRA MIR (Absorption)
Print this report.
ManuelMarchio - 11:57, Wednesday 17 June 2015 (107)Get code to link to this report
AC noise -Lockin Time constant; Photodetector noise; Pump power needed

I acquired 3 minutes of the AC noise in the following conditions;

 - pump off

 - probe on

 - front door closed

  - various lock in integration time constants;

As shown in the plot , the higher the time constant, the lower the noise.

But we have to choose it depending on the total acquisition time, for example we cannot spend a week to acquire a surface map.

I report the values:

timeconstant (ms)

Noise std (microV)
10 47
30 14
100 18
300 9.5
1000 5.2

We remind the specification of the system says between 5 and 25 microV at time constant: 100ms; and chopper frequency: 380Hz.

-----------------------------------------------------------------------------------------------------------------------------------------------------------------------------

Then we switch off the HeNe probe and measure the noise of the photodetector with 100ms time constant.

The photodetector noise standard deviation is 0.8 microV (see the histogram below) and it includes the detector noise and the lockin noise

-----------------------------------------------------------------------------------------------------------------------------------------------------------------------------

For absorption of 22% we have about 350mV AC signal at 30mW pump power;

The signal should be at least 2 or 3 times the noise (18microV * 3);

For 12ppm we expect 19microV of AC signal at the same pump power (30mW)

This means that a pump power of 100mW might be sufficient to see a good signal for 12ppm. It follows that we need 10W to see absorptions of 0.1ppm

Images attached to this report
107_20150617035217_noisestdvstimeconst.jpg 107_20150617041842_probeoffhisto.jpg
KAGRA MIR (Absorption)
Print this report.
ManuelMarchio - 10:36, Wednesday 17 June 2015 (106)Get code to link to this report
AC noise - Chopper frequency

I tuned the lockin gain to have a better resolution for little signals.

I acquired 3 minutes of noise for different modulation frequencies, in the following conditions:

 - Probe ON;

 - Pump OFF;

 - Front door closed;

 - Lockin time constant 100ms;

 

I made an histogram for every acquisition and I computed the standard deviation in order to find the best chopper frequency.

The lowet value of the noise standard deviation is at 430 Hz modulation frequency. The second best value is for 480 Hz, as used previously. The highest noises are at multiples of 50 Hz, as expected.

Images attached to this report
106_20150617033133_histograms.jpg 106_20150617033153_stdnoisechopper.jpg
KAGRA AOS (General)
Print this report.
TomotadaAkutsu - 18:51, Tuesday 16 June 2015 (102)Get code to link to this report
Oplev's pillar test installation to the tunnel

Abstract: we did a test installation of an oplev's pillar to the tunnel.

Workers: Uchiyama and Akutsu

The test installation has been done around (1) PR3 chamber, (2) IFI chamber (for IMMT2 mirror), and (3) MCF chamber (for MCo mirror).

For all cases, the bolt holes on the floor are all dirty, and the M24x130 bolts are hard to fit to them without cleaning the bolt holes; even after the cleaning, it was so hard,

so ratchet socket wrenches are used to screw. Actually we didn't have M24 washers today, so the legs have not been fully fixed.

 (1) for PR3,

It is known that the pillar and the chamber's hinge will mechanically interfer, and the inconvenient fact was confirmed. We just slide the pillar a little from the planned location, and found a so-so place, but not know wether the optical table on the pillar can be installed or not. When the pillar stands, the top surface is almost horizontal without shims (shown by a whater leveler), and this means the floor surface here and the pillar's upper and lower surfaces are well manufactured.

(2) for IMMT2 mirror

Not much mechanical intereferences, but one concern is space between a barrelhead attached to a gate valve between MCF and IFI chambers. Still the bolts were hard. When the pillar stands, the top surface is almost horizontal without shims (shown by a whater leveler), and this means the floor surface here and the pillar's upper and lower surfaces are well manufactured. The height from the top surface to the beam line is measured to be 112.5 mm (the beam line height is just estimated from the chamber structure). Then the oplev beam height from the optical table on the pillar can be estimated 112.5 - 27 = 85.5 mm. The expected beam height was 85.85~85.9 mm, so it could be in the tolerant range.

(3) for MCo mirror

Not much mechanical intereferece, but one concern is space between a reducing flange and the pillar. Still the bolts were hard. Unfortunately, the floor surface for the pillar is not so good, and we need two pieces of shims to make the top surface horizontal. The height from the top surface to the beam line is measured to be 115 mm (the beam line height is just estimated from the chamber structure), which should be the same as the one arund IFI chamber, but now it is obvious to have 2.5 mm difference. This difference would be a little bit large if you take into consderaton about measurement errors. Then the oplev beam height from the optical table on the pillar can be estimated 115 - 27 = 88 mm. The expected beam height was 85.85~85.9 mm, so it could be a little bit larger than my expectation, but can be tuned.

Images attached to this report
102_20150616114556_dcim1209.jpg 102_20150616114613_dcim1211.jpg 102_20150616114703_dcim1215.jpg 102_20150616114722_dcim1217.jpg 102_20150616114807_dcim1222.jpg 102_20150616114825_dcim1224.jpg 102_20150616114953_dcim1226.jpg 102_20150616115049_dcim1227.jpg 102_20150616115107_dcim1237.jpg 102_20150616115116_dcim1235.jpg
KAGRA AOS (General)
Print this report.
SimonZeidler - 17:33, Tuesday 16 June 2015 (104)Get code to link to this report
PSD calculations

I got the two-dimensional map of the BS from Hirose-san.
I will try to implement these data into the "LightTools" software in the next days.

At the same time, I tried to fit the one-dimensional PSD with the K-correlation model with only limited success.
It seems that this particular model is not suitable for the data...or the data aren't enough.
From the map it should also be possible to calculate the two-dimensional PSD and to fit these data instead of the one-dimensional one.
I would like to try these calculations also during the next days to see how far I can use the PSD for the scattering Paper I am now writing.

KAGRA AOS (General)
Print this report.
SimonZeidler - 17:27, Tuesday 16 June 2015 (103)Get code to link to this report
Updated JASMINE scatterometer

The laser for the JASMINE scatterometer is now focussed on the sample.
I used a lens with f=200mm; the size of the focal point is now 0.25 - 0.3 mm.

Measurements can again be taken.

I will do some test measurements tomorrow on a Titanium sample.

KAGRA VIS (OSEM)
Print this report.
TomotadaAkutsu - 20:16, Monday 15 June 2015 (101)Get code to link to this report
Flexi circuits

Finally Tatsumi-san gets the (usable) flexi circuit prototypes for OSEMs, and he gave them to me. I firstly check whether the soldering can be done, and the results so far are good.

Images attached to this report
101_20150615131621_dcim1179.jpg 101_20150615131637_dcim1182.jpg
KAGRA AOS (General)
Print this report.
TomotadaAkutsu - 18:44, Friday 12 June 2015 (100)Get code to link to this report
Assy of Oplev's box prototype

An oplev's covering box prototype is assembled. Needs more additive modifications.

Images attached to this report
100_20150612114602_oplevboxs.jpg
KAGRA MIR (Absorption)
Print this report.
ManuelMarchio - 17:54, Friday 12 June 2015 (99)Get code to link to this report
AC noise

 I measured the AC signal to estimate the noise in the following conditions:

No sample installed ;

Chopper frequency: 478 Hz ;

Lock-in time costant: 100 ms ;

Tuned to maximum DC signal ;

Front door closed ;

Motors moving but time wait 500 ms ;

 

Measure are taken in 3 different situations

 - Pump open at 30mW --> figure empty20150612.jpg ;

 - Pump  closed --> figure emptyclosed20150612.jpg ;

 - My voice singing around 478 Hz like here --> figure manuelvoice478hz.jpg ;

I measured 60 microV with silent, and 2.2 mV with acustic noise at chopper frequency.

 

The manual  says, at 380 Hz and with motors not moving, the noise is between 5 and 25 microV.

I will make measurement for many chopper frequencies and look for the minimum.

Images attached to this report
99_20150612102939_empty20150612.jpg 99_20150612102948_emptyclosed20150612.jpg 99_20150612103006_manuelvoice478hz.jpg
KAGRA MIR (Absorption)
Print this report.
DaisukeTatsumi - 05:28, Thursday 11 June 2015 (97)Get code to link to this report
Spectrophotometry of two reference samples

By using spectrophotometer (Shimadzu SolidSpec 3700) at NAOJ ATC

transmissivity and reflectivity of the reference samples were measured.

From these measurements, absorption value can be obtained.

SURFACE SAMPLE:    Newport FRQ-ND02          22.0%

BULK SAMPLE:            Schott glass  NG-12         37.5%     ---> 37.5% / 0.36 cm = 104% /cm

Both of values are almost consistent with catalog spec. (22.2% and 116% /cm).

Images attached to this report
97_20150611022329_bulkreferencertl.png 97_20150611022336_surfacereferencertl.png
Non-image files attached to this report
KAGRA VIS (Mirror Assemble)
Print this report.
DaisukeTatsumi - 10:19, Wednesday 10 June 2015 (96)Get code to link to this report
Glueing test of MasterBond EP30-2

I did the following things.

----------------------------------------------------------------------------------------------------------------------

[ Mixuture ]

EP30-2 (A) : EP30-2 (B) : borosilica = 10 : 1 : 0.55 = 5 g : 0.5 g : 0.275 g

-----------------------------------------------------------------------------------------------------------------------

[ Cure schedule ]

24 - 48 hours        at 75 Fdeg = 21 Cdeg

2.0 - 2.5 hours      at 200 Fdeg = 93 Cdeg

-----------------------------------------------------------------------------------------------------------------------

Test 1: Cure Testing

Provide heating by hot plate. Set the temperature to be 90 Cdeg (~200 Fdeg).

Leave the glue for 15 minutes. After the heating, cool it for 1 minute.

Because it is smooth and hard, it is a well-mixed batch.

------------------------------------------------------------------------------------------------------------------------

Test 2: Glue flag and flag base

After the test 1, I made glueing two sets of (dummy) flag and flag base.

One is heated at 90 Cdeg for 3.5 hours. Another is placed at room temperature (~24 Cdeg).

-----------------------------------------------------------------------------------------------------------------------

Images attached to this report
96_20150610041517_mixtureofglue.jpg 96_20150610041608_curetesting.jpg 96_20150610041701_curedglue90cdeg15minp2.jpg
KAGRA AOS (General)
Print this report.
TomotadaAkutsu - 18:34, Tuesday 09 June 2015 (95)Get code to link to this report
iKAGRA AS port BRT design

Uploaded a draft of the design document for the iKAGRA AS port BRT (in Japanese)

http://gwdoc.icrr.u-tokyo.ac.jp/cgi-bin/private/DocDB/ShowDocument?docid=3624

I and Michimura-kun also had a skype meeting on how to take care this task...

KAGRA VIS (Mirror Assemble)
Print this report.
DaisukeTatsumi - 12:48, Tuesday 09 June 2015 (94)Get code to link to this report
Flag thickness measurement

To design the ASSEMBLE JIG,

I made measurement of flag thickness as attached.

Non-image files attached to this report
KAGRA VIS (OSEM)
Print this report.
TomotadaAkutsu, KokiOkutomi - 21:04, Monday 08 June 2015 (93)Get code to link to this report
OSEM preliminary test with a new test bench

OSEMs are preliminarily tested with a new OSEM test bench in a non-clean environment. The OSEM's LED and PD are operated by a test circuit (not an actual satellite box, but I and Okutomi-kun have just copied requred parts...)

The conbinations under test are:

  1. New LED holder with TSTS7100 and new PD holder
  2. Old LED holder with OP232 and new PD holder (NOTE!)
  3. Old LED holder with OP232 and old PD holder (NOTE)

NOTE: the "old" LED holder cannot hold a lens with the same manner as the ones used in TAMA300 SAS prototype now, as it is revealed that the adhesive of lens to the "lens holder" inside the LED holder cannot be usable in vacuum. I just make the lens holded somehow....... so the distance from the LED and the lens are changed. However, from the rough measurement of beam profile of light of the ones, I got impression that such a change of positions doesn't have effects on the performance... (preliminary)

What to be confirmed is the linear range. I'd like to compare the range with the one used in TAMA300 SAS prototype.

So far the linear range are almost all the same.

Obviously when the linear range is such, we don't need to use such a large area photodiode; it is too large and hard to fit new and old designed PD holder. A small aperture PD would have smaller capacitance, and that makes the cricuit easliy stabilized.

... and 4. Plot of data by Sekiguchi-kun for TAMA-OSEMs; the data's range differs from our setup so I just lineary scaled the vertical axis. It appears that the linear range would not change.

Images attached to this report
93_20150608135449_dcim1158.jpg 93_20150609134424_figure1.png
KAGRA AOS (General)
Print this report.
SimonZeidler - 17:14, Monday 08 June 2015 (92)Get code to link to this report
Outgassing results of VB and MV

Regarding the Doughnut Baffle, the two promising candidates for a very good absorption/scattering coating have been tested for their ultra-high vacuum compatibility.

The results show that VB has better UHV capabilities then MV. Especially water seems to be highly abundant on the surface of MV.

Images attached to this report
92_20150608101356_1.jpg
KAGRA VIS (General)
Print this report.
RyutaroTakahashi - 15:43, Monday 08 June 2015 (91)Get code to link to this report
Replacement of the breadboard in MCE
Member: N. Ohishi, H. Ishizaki, M. Barton, R. Takahashi

May 26
Checking condition of the clean booth at 1530-17:00 by Takahashi.
1) Turned on all (16) filters. 4 filters had been working before.
2) Hirose-san measured particles by the large counter. I measured particles by the small counter too. Number of particles was 5/L just after turnung on the filters. Humidity was 80.5%.
3) Hirose-san measured particles continuously during I walk around.

May 27
Replacement of the breadboard (1).
1) Cleaned the floor of the clean booth by the vacuum cleaner and the mope.
2) Opened the chamber.
3) Moved the optical bench into the clean boooth by the jib crane.
4) Extracted the suspension from the chamber and put it on the optical bench.
5) Extracted the breadboard. Up-down movement of the crane (chain block) is very slow.
6) Rapped the breadboard.

May 28
Replacement of the breadboard (2)
1) Cleaned the floor of the clean booth by the mope.
2) Cleaned inside the chamber by the vacuum cleaner, the wet wiper, and the dry cloth.
3) The SUS blocks were stacked after first disassembling. Since the guid rods are inside the MCF chamber, we don't use them. The crane can not move to ransverse direction. This is a problem.
4) Unpacked the renewaled breadboard and installed it into the chamber.
5) The breadboard extracted was enclosed by the wood box.
6) Rapped the suspension with decicant.
7) Closed the chamber.
KAGRA AOS (General)
Print this report.
TomotadaAkutsu - 19:48, Friday 05 June 2015 (90)Get code to link to this report
Oplev Reciever

Now a prototype of the oplev receiver

http://gwdoc.icrr.u-tokyo.ac.jp/cgi-bin/private/DocDB/ShowDocument?docid=3562

is well assembled. Note that the QPD circuit in the photograph below has a wrong connector, and should be modified soon.

Images attached to this report
90_20150605124759_dcim1153.jpg
KAGRA VIS (General)
Print this report.
NaokoOhishi - 15:30, Thursday 04 June 2015 (89)Get code to link to this report
MCE Breadboard Replacement

27th-28th May,

Takahashi-san, Mark-san, Ishizaki-san, and Ohishi changed MCE breadboard.

Due to the unexpected stack cleaning work, surface level adjustment of the MCE board was postponed to 11th-12th Jun.

Inside the cleanbooth, humidity was  - 80% but cleanness was good -ISO class 4.

Optical bench was carried into the booth.

Attached pictures were taken by Koichiro (an Olympus camera equipped to NAOJ Kamioka branch).

Images attached to this report
89_20150604080444_p5270049.jpg 89_20150604080802_p5280081.jpg 89_20150604080820_p5280097.jpg 89_20150604080939_p5280082.jpg 89_20150604081033_p5280090.jpg 89_20150604082335_p5280074.jpg
KAGRA MIR (Absorption)
Print this report.
DaisukeTatsumi - 11:22, Thursday 04 June 2015 (88)Get code to link to this report
Calibration with standard samples

[SURFACE SAMPLE]

Newport FRQ-ND02

Absorption   22.2%  at 1064 nm

pump power damage threshold 40mW

------------------------------------------------------------------------------------------------------------------------------------------------

Type UV Fused Silica Metallic ND Filter   ---> Inconel coated fused silica substrate

Optical Density                    0.2 at 633 nm 

Diameter                              1.0 in. (25.4 mm)

Thickness                            0.118 in. ( 3.0 mm)

Material                                UV grade fused silica

Surface Quality                    60-40 scratch-dig

Surface Flatness                  less than 1 lambda at 632.8 nm over the clear aperture

Wedge                                  less than 3 arc min

Wavelength Range              350 - 2000 nm

Diameter Tolerance             ±0.25 mm

Thickness Tolerance            ±0.25 mm

Clear Aperture                      central 80% of dimensions

Damage Threshold               30 W/cm2 CW, typical;

                                              not recommended for pulsed laser use

Angle of Incidence                0 degree

Transmission                        63.1 % at 632.8 nm

Optical Density Tolerance    ±7%

Chamfers                              0.25-0.76 mm face width

Chamfers Angle/Tolerance   45 ±15 degree

Cleaning                                Non-abrasive method, acetone or isopropyl alcohol on lens tissue recommended

[BULK SAMPLE]

Schott glass NG-12

Thickness                                      3.6 mm

Absorption rate                              116% /cm

Refractive index                             1.48 at 1014 nm

Pump power damage threshold     100 mW

--------------------------------------------------------------------------------------------------------

Absorption = 1.16 /cm * 0.36 cm = 0.418 = 41.8%

Images attached to this report
88_20150604035828_bulksample0150602.jpg 88_20150604035846_20150602surface.jpg
KAGRA VIS (Mirror Assemble)
Print this report.
DaisukeTatsumi - 10:42, Wednesday 03 June 2015 (87)Get code to link to this report
Dummy flag for glueing test

Dummy flag and (prototype) flag base were delvered as attached file.

By using these parts I will try to adhere.

Non-image files attached to this report